# Duplex-GS: Proxy-Guided Weighted Blending for Real-Time Order-Independent Gaussian Splatting

### v1

@@ III-C Weighted Sum Rendering
By employing this WSR formulation, the non-commutative, computationally intensive sorting stage—traditionally required for compositing—is eliminated. This significantly enhances the potential for parallelization and accelerates rendering. Specifically, the SOTA Linear Correction weighted-sum rendering (LC-WSR) function $w(d_{i})$ is defined as
@@
IV Layer I: Cell Rasterization for 3DGS
While neural Gaussian frameworks [16, 17] have achieved remarkable reconstruction accuracy, the use of anchors as encoding proxies lacks a concrete geometric interpretation. So we redesign the neural Gaussian proxies to alleviate the computational overhead associated with sorting and rasterization. Specifically, instead of encoding 3D Gaussians with anchors, we introduce the concept of “cell”—an ellipsoid with explicit geometric properties and a defined visible region—to encode neural Gaussians together with their latent features. Based on this representation, we further propose a cell rasterization strategy, enabling efficient and lightweight rasterization.
@@ IV-A Cell Proxy
Unlike anchors, which are points lacking geometric structure, each cell is represented as an ellipsoid that contains several neural 3D Gaussians whose physical properties are dynamically decoded from a trainable feature.
Initialization begins with SfM points $\mathbf{P}\in\mathbb{R}^{N\times 3}$, from which we generate the initial cells, each centered at a point in accordance with the method described in [17]. Every cell is parameterized by physical structure scales $\bm{S}_{n}$ and and quaternions $\bm{Q}_{n}$, defining its visible region. In the dual-layered hierarchy, each cell is further associated with a latent feature vector $\bm{f}_{n}$ that encodes local scene information, and emits $K$ neural Gaussians as described in Eq. 4. The attributes of these $K$ neural Gaussians—including color $\bm{c}_{n,k}$, scales $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$—are decoded from the latent feature $\bm{f}_{n}$, along with the viewing distance $\delta_{n}$ and viewing direction $\bm{d}_{n}$, via shared scene-wise MLPs.
To ensure that each Gaussian center lies within the visible region of its corresponding cell, we constrain the offset parameter $\mathcal{O}_{n,k}^{\text{pos}}$ to the range [1,1][-1,1]. Additionally, to align the collective regions of the emitted neural Gaussians with the visible extent of the cell, we modify the encoding of the scale $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$ attributes, thereby directly linking the geometric properties of the cell and its Gaussians
@@
$\{\bm{q}_{n,1},...,\bm{q}_{n,K}\}=F_{q}(f_{n},\delta_{n},d_{n}),$ (10)
where the output of $F_{s}$ and $F_{q}$ are activated by sigmoid and normalization respectively. Furthermore, the overall geometric configuration of each cell is explicitly associated with that of its constituent Gaussians through
@@
$\bm{Q}_{n}=\text{norm}(\sum_{k}^{K}\bm{q}_{n,k}*\sigma_{n,k}),$ (12)
To restrict that the Gaussian regions fully locate inside the cell region, the maximum of $\mathcal{O}_{n,k}^{\text{scale}}$ is clamped by
@@
IV-B Cell Searching Rasterization
In the original 3DGS framework [7], views are rendered by rasterizing 3D Gaussians onto the image plane through a Gaussian search rasterization strategy. In contrast, the anchors introduced in [16], which form the proxies of the dual-layered hierarchy, do not directly participate in the rendering process; instead, rendering remains confined strictly to the Gaussian level. This results in a distribution mismatch between the two hierarchy levels, i.e., between anchors and 3D Gaussians.
Building on our proposed geometric cell, we introduce a cell search rasterization technique in which cells, rather than individual Gaussians, serve as the primary rasterization units. As illustrated in Fig. 3, we perform rasterization at the cell level during the pre-processing stage, which reduces memory access overhead. Specifically, each cell defines a visible region, and only those whose regions intersect with the viewing frustum are recorded and subsequently decoded into neural 3D Gaussians for the subsequent $\alpha$-blending stage. This strategy ensures that the distribution of cells and 3D Gaussians are naturally aligned across the two hierarchy levels, and restricts rasterization exclusively to the cell level. Following the tile based paradigm, cells are assigned to individual tiles, and radix sort is applied to the cell sequence as necessary. Because each cell encapsulates a local cluster of Gaussians, the total number of cells is substantially less than the number of individual Gaussians. As a result, both memory consumption and computational overhead during rasterization are significantly reduced.
Figure 4: Comparison of three different transmittance: LC-WSR (Eq. 8), vanilla 3DGS (Eq. 3) and ours (Eq. 16). The proposed transmittance is not strictly monotonically decreasing with increasing depth, indicating that the Gaussians are locally unordered. Nevertheless, our method achieves blending weights similar to those of vanilla 3DGS, without requiring an explicit Gaussian order.
@@ V Layer II: Physical Weighted Sum Rendering
Duplex-GS reconstructs 3D scenes using ellipsoidal cells, enabling rasterization and sorting to be performed at the cell level, thereby significantly reducing rasterization overhead. However, without explicitly sorting the depths of individual Gaussians—a requirement in the original $\alpha$-blending operations grounded in physical modeling—the exact blending order becomes ambiguous. Sort-free Gaussians [12], an OIT extension of vanilla 3DGS, demonstrated that optimized components and parameters for high-fidelity view synthesis are governed solely by the mathematical constraints of the scene model. This weighted sum approach is well-suited for integrating 3DGS into modern GPU graphics pipelines and for deployment on resource-constrained devices. Nevertheless, abandoning physical constraints introduces “transparency” artifacts that degrade visual quality, and the absence of early termination further results in reduced rendering efficiency.
@@ V-A Physical OIT Blending for 3DGS
WSR [12] is a mathematically motivated rendering model, in which $\alpha_{i}$ in Eq. 6 does not directly represent the meaningful opacity of a Gaussian and may even exceed one due to its lack of a physical basis. While this formulation facilitates the weighted sum rendering process, it inherently precludes the early-termination mechanisms used in conventional $\alpha$-blending.
Obtaining the sorted order at the cell level is both memory and computationally efficient, since the number of cells is substantially fewer than that of the 3D Gaussians. Leveraging this, we redefine $\alpha$ as the opacity weighted by the Gaussian kernel, as specified in Eq. 2, constraining its value to the physically meaningful range of $[0,1]$.
Thanks to the approximate ordering provided by cell-level sorting—which adequately reflects the relative relationships among the contained Gaussians—we are able to enforce physical constraints within the mathematical WSR framework, without the need for precise ordering of all Gaussians. As illustrated in Fig. 3, the final rendered view $\bm{C}$ is computed according to the weighted sum rendering formula:
@@
where $T_{n}^{\text{cell}}$ denotes the cell-level transmittance, which is efficiently obtained by sorting the sparsely distributed cells and can be calculated as:
$T_{n}^{\text{cell}}=\prod_{i=1}^{n-1}\prod_{k=1}^{K}(1-v_{i}\cdot\alpha_{i,k}).$ (16)
Due to the non-uniform distribution and potential overlap of cells, $T_{i}^{\text{3DGS}}$ in Eq. 3 may not precisely capture the transmittance of all emitted neural Gaussians. The introduction of the learnable parameter $v_{n}$ for each cell effectively compensates for this approximation. Three relevant paradigms of transmittance and the resulting Gaussians weights are visualized in Fig. 4. LC-WSR simulates the decay trend of transmittance with a linear function w.r.t. Gaussian depth, introducing great number of Gaussians with tiny weights. Our method obtain similar weights compared with the vanilla 3DGS without the requirements of concrete order of individual Gaussians.
@@ V-B Early Termination for Cell Based 3DGS
We propose to perform sorting based on cells, rather than individual Gaussians, thereby significantly reducing sorting overhead while retaining the hardware-friendly weighted sum rendering paradigm.
As $T_{n}^{\text{cell}}$ decreases during the front-to-back rendering of cells, early termination can be efficiently incorporated to accelerate the rendering process and mitigate transparency artifacts. Specifically, $T_{n}^{\text{cell}}$, defined in Eq. 16, approaches zero as cell-level transmittance accumulates and saturates, which provides a practical criterion for early termination: the blending process is halted when $T_{n}^{\text{cell}}<\epsilon$. Furthermore, in contrast to previous WSR methods for 3DGS, our approach restores explicit physical meaning to the opacity parameter $\sigma_{n,k}$, restricting it to the range $[0,1]$. This property facilitates pruning and regularization during training.
@@ VI Optimization Strategies for Cell Based Gaussian Splatting
The proposed cell primitives are endowed with distinct physical attributes; however, these features are not directly involved in the differentiable blending process. Furthermore, advanced optimization strategies are introduced to systematically organize the evolving cell structures.
VI-A Alignment of Cell and Gaussian Region
The cell search rasterization inevitably leads to the introduction of redundant 3D Gaussians during the blending stage, due to the discrepancies that arise between the spatial intervals of cells and the emitted neural Gaussians. Furthermore, the efficiency of cell search rasterization is contingent upon the degree of coherence between the visible regions of a cell and its associated Gaussians.
VI-A1 False positive skip
@@
TBR is widely adopted in modern GPUs and is also utilized in 3DGS rasterization. However, this approach often introduces a substantial number of redundant false-positive Gaussians, a problem that is further exacerbated in cell based Gaussian frameworks due to the spatial mismatch between cell regions and the actual coverage of Gaussians. To mitigate the associated performance degradation, we introduce a pixel-level radius based culling strategy to eliminate unnecessary Gaussian kernel computations. Specifically, as illustrated in Fig. 5, a lightweight axis-aligned bounding box (AABB) check is employed: Gaussians are excluded from the blending operation if the pixel lies outside their respective AABBs. This efficient early-exit mechanism significantly reduces redundant computation and accelerates the overall rendering process.
VI-A2 Geometric correction via neural Gaussians
In addition to skipping false-positive Gaussians to improve efficiency, maximizing the overlap between each cell and its emitted neural Gaussians further reduces the number of redundant Gaussians. To achieve better geometric alignment, we dynamically define the cell’s quaternion as a weighted average of the quaternions associated with its emitted Gaussians, as described in Eq. 12. Moreover, to ensure an even spatial distribution of Gaussians within each cell, the cell centers are periodically updated according to
@@ VI-A2 Geometric correction via neural Gaussians
VI-A3 Cell position reset
As discussed in Gaussian based models [7, 16, 17], the initial SfM point cloud often fails to comprehensively cover the entire scene, particularly in textureless or sparsely observed regions. Following the adaptive density control (ADC) policies employed in previous works [7], we utilize view-space positional gradients as a densification score to guide the growth of new cells. Specifically, additional cells are introduced at the centers of Gaussians that attain high densification scores. In contrast to the anchors proposed in [16, 17], which do not participate directly in the rendering process, our cells possess explicit physical attributes and independently encode local regions. To decouple existing cells from newly introduced ones, the offsets of the significant Gaussians are reset to zero after cell growth, i.e., their positions are reinitialized to the centers of the corresponding cells.
VI-B Perceptual Loss
The loss function for optimizing 3D Gaussian Splatting (3DGS) is typically formulated as a combination of $\ell_{1}$ and SSIM [33] losses, which respectively capture pixel-level reconstruction error and structural similarity. Minimizing these terms promotes high signal-to-noise ratios but often results in images that lack high-frequency details and appear perceptually unsatisfying, particularly at higher resolutions. Given the effectiveness of perceptual loss functions in image generative models [34, 35], we additionally incorporate a perceptual loss term for training Gaussian based models. The overall loss is defined as
@@ VII Experimental Results
    The official implementation has not been released. Experiments are conducted using our own implementation, which is available at https://github.com/LiYukeee/sort-free-gs.
@@
    The official implementation has not been released. Experiments are conducted using our own implementation, which is available at https://github.com/LiYukeee/sort-free-gs.

### v2

@@ III-A2 Weighted Sum Rendering
By employing this WSR formulation, the non-commutative, computationally intensive sorting stage, traditionally required for compositing, is eliminated. This significantly enhances the potential for parallelization and accelerates rendering. Specifically, the Linear Correction Weighted-Sum Rendering (LC-WSR) function $w(d_{i})$ achieve best performance which is defined as
@@
III-B Overview
We introduce a novel framework featuring a tailored encoding and rendering paradigm, as illustrated in Fig. 3. Our approach is structured in two primary stages. First, 3D scenes are modeled using ellipsoidal cells, which act as proxies that emit Gaussians with coherent features inside their coverage. This is coupled with a hybrid rendering paradigm that integrates the advantages of $\alpha$-blending and WSR, establishing a globally ordered yet locally unordered scene structure. Specifically, rasterization is first performed at the proxy layer to compute a blending weight for each cell in $\alpha$-blending manner. The emitted Gaussians are then rendered using a revised WSR paradigm.
This pipeline conducts sorting at the cell level, establishing a coarse spatial order for patches of Gaussians. This strategy drastically reduces the computational burden of performing a radix sort on every individual Gaussian primitive (Sec. IV). Within each cell, Gaussians remain locally unsorted, making the order-independent WSR paradigm naturally suitable for their blending. This design eliminates popping artifacts, as the rendering result is locally invariant to the Gaussian order. Furthermore, by reintroducing physical constraints into the OIT-based renderer, a physically grounded transmittance decay is guaranteed, as shown in Fig. 5, which removes the transparency artifacts inherent in the vanilla OIT-based Gaussian Splatting [16] (Sec. V) as well as enables early termination in cell-level to accelerate rendering.
IV Layer I: Cell Proxy for 3DGS
While neural Gaussian frameworks [20, 21] have achieved remarkable reconstruction accuracy, they still rely on conventional $\alpha$-blending operations as the introduced proxy anchors are points without concrete geometric interpretation. We introduce ellipsoidal cells with explicit visible regions as the proxies. Based on this representation, a cell rasterization strategy is introduced, enabling efficient and lightweight rasterization.
Figure 4: Comparison between rasterization of GS and cell. Unlike the original $\alpha$-blending which processes each Gaussian individually, our method performs rasterization and early termination at the coarser cell level, where each cell encapsulates multiple Gaussians. The order confusion caused by overlap among cells is discussed in Sec. V-A.
@@ IV-A Cell Proxy
Unlike anchors, which are points lacking geometric structure, each cell is represented as an ellipsoid that covers several neural 3D Gaussians whose physical properties are dynamically decoded from a trainable feature.
Initialization begins with SfM points $\mathbf{P}\in\mathbb{R}^{N\times 3}$. We generate the initial cells from these points, centering each one in accordance with the method described in [21]. Every cell is parameterized by position $\bm{x}_{n}$, scales $\bm{S}_{n}$, quaternions $\bm{Q}_{n}$, defining its visible region, and a calibration scalar $v_{n}$ (discussed in Sec. V-A). In the dual-layered hierarchy, each cell is further associated with a latent feature vector $\bm{f}_{n}$ that encodes local scene information, and emits $K$ neural Gaussians as described in Eq. 1. The attributes of these $K$ neural Gaussians, including color $\bm{c}_{n,k}$, scales ratio $\mathcal{O}_{n,k}^{\text{scale}}$ and quaternions $\bm{q}_{n,k}$, are decoded from the latent feature $\bm{f}_{n}$, along with the viewing distance $\delta_{n}$ and viewing direction $\bm{d}_{n}$, via shared scene-wise MLPs.
$\{\bm{c}^{n,1},...,\bm{c}^{n,K}\}=F_{c}(f_{n},\delta_{n},d_{n}),$ (8)
@@
$\{\bm{q}_{n,1},...,\bm{q}_{n,K}\}=F_{q}(f_{n},\delta_{n},d_{n}).$ (10)
To ensure that each Gaussian center lies within the visible region of its corresponding cell, we constrain the offset parameter $\mathcal{O}_{n,k}^{\text{pos}}$ to the range [1,1][-1,1]. We propose that cells are likely to emit Gaussians with similar geometric properties. So scales of Gaussians are derived under the supervision of $\bm{S}_{n}$:
@@
where the output of $F_{s}$ is activated by sigmoid to restrict the ratio $\mathcal{O}_{n,k}^{\text{scale}}$ to the range $(0,1)$. Based on Eq.11, the scales of cells can be directly optimized. While the quaternions unable to optimize in the similar way, as they denote the orientation of the clustered Gaussians. To align regions of cell and Gaussians, the cell quaternions are updated as the weighted average of the associated Gaussians’ orientation
$\bm{Q}_{n}=\text{norm}(\sum_{k}^{K}\bm{q}_{n,k}*\sigma_{n,k}).$ (12)
Additionally, to restrict that the Gaussian regions fully locate inside the cell region, the maximum of $\mathcal{O}_{n,k}^{\text{scale}}$ is further clamped by
@@
IV-B Cell Rasterization
Conventional $\alpha$-blending renders views by rasterizing 3D Gaussians onto the image plane via splatting, a process that necessitates a global Gaussian traversal and sort. We fundamentally reformulate this pipeline by introducing a cell-level rasterization technique, where ellipsoidal cells, rather than individual Gaussians, serve as the primary rendering primitives. As illustrated in Fig. 4, this approach significantly reduces the number of primitives handled. Specifically, each cell defines a bounded visible region; only cells intersecting the viewing frustum are selected and decoded into neural 3D Gaussians.
This proxy-guided design inherently ensures spatial alignment between the cell distribution and the underlying Gaussians, confining computationally intensive rasterization operations exclusively to the coarser cell level. Operating within a tile-based paradigm, visible cells are assigned to corresponding tiles, and a radix sort is applied to the cell sequence. Since the number of cells is orders of magnitude smaller than that of Gaussians, this strategy yields substantial reductions in both memory consumption and computational overhead for the rasterization process.
Figure 5: Comparison of three different transmittance: LC-WSR (Eq. 7), vanilla $\alpha$-blending (Eq. 4) and ours (Eq. 16). By reintroducing physical constraints, our model produces a transmittance curve similar to that of $\alpha$-blending, establishing it as a reliable criterion for early termination. Furthermore, the proposed transmittance is not strictly monotonic with depth, confirming the local order-independent nature of Gaussian blending in our framework.
@@ V Layer II: Physical Weighted Sum Rendering
Duplex-GS reconstructs 3D scenes using ellipsoidal cells, which enable rasterization and sorting at a coarse cell level to significantly reduce rasterization overhead. However, this approach forgoes the explicit per-Gaussian depth sorting required by the original $\alpha$-blending pipeline, which is grounded in physical light transport models. This introduces an ambiguity in the blending order. The Sort-free Gaussian method [16], an OIT extension of 3DGS, demonstrates that high-fidelity view synthesis can be achieved through a weighted sum of Gaussians, governed by mathematical rather than physical constraints. This paradigm is highly suitable for graphics rendering pipelines in modern GPU and resource-constrained devices. Nevertheless, abandoning the physical blending model introduces transparency artifacts that degrade visual quality, and the lack of early termination further compromises rendering efficiency.
@@ V-A Physical OIT Blending for 3DGS
Although the WSR paradigm [16] provides a mathematically-grounded rendering model, it lacks a physical foundation. Consequently, the blending weight $\alpha_{i}$ in Eq. 5 does not correspond to a physically meaningful opacity and is unconstrained, often exceeding unity. This formulation not only introduces transparency artifacts but also precludes the use of early termination when blending opaque surfaces, limiting its efficiency.
Within this coarse-ordered structure obtained from cell rasterization, we redefine $\alpha$ as an opacity weighted by the Gaussian kernel (Eq. 3), constraining its value to the physically plausible range of $[0,1]$. The final rendered view $\bm{C}$ is computed according to the weighted sum rendering formula:
@@
where $T_{n}^{\text{cell}}$ denotes the cell-level transmittance, which is efficiently obtained by sorting the sparsely distributed cells, and it decays based on the accumulated opacity from preceding cells:
$T_{n}^{\text{cell}}=\prod_{i=1}^{n-1}\prod_{k=1}^{K}(1-v_{i}\cdot\alpha_{i,k}),$ (16)
where a learnable calibration scalar $v_{i}$ is introduced to address the ambiguity arising from non-uniform spatial distribution and potential inter-cell overlaps, as illustrated in Fig. 4. Since cell occlusion relationships are view-dependent, we parameterize $v_{i}$ using spherical harmonic coefficients to capture its directional variation. The resulting Eq. 14 shows that our framework produces model which is ordered outside cells and order-independent inside each cell.
The opacity distribution in a 3D scene is typically sparse, implying that the accumulated transmittance is physically modeled by a step-function which suddenly changes encountering points with high opacity rather than a gradual decay. To validate this theory, we provide three transmittance paradigms and the resulting Gaussians weights in Fig. 5. LC-WSR simulates transmittance with a linear function of Gaussian depth. However, this model is physically inconsistent, as it is derived from purely mathematical constraints. In contrast, our method produces a weight distribution similar to vanilla $\alpha$-blending without requiring a precise per-Gaussian depth order, which eliminates transparency artifacts for that the transmittance decays rapidly when encountered opaque object. Furthermore, the rendering of Gaussian primitives is locally order-independent, as the transmittance does not strictly decrease with depth. This property forms the cornerstone for the elimination of popping artifacts.
@@ V-B Early Termination for Cell Based 3DGS
As defined in Eq. 16, the cell-level transmittance $T_{n}^{\text{cell}}$ decays towards zero during front-to-back rendering, eliminating transparency artifacts. This provides a natural mechanism for early termination: the blending process is halted in cell level once $T_{n}^{\text{cell}}$ falls below a predefined threshold $\epsilon$, as illustrated in Fig. 4. This approach effectively accelerates rendering. Unlike prior WSR methods for 3DGS, our formulation restores explicit physical meaning to the opacity parameter $\sigma_{n,k}$ by constraining it to the physically plausible range of $[0,1]$, which means the empirical threshold can be directly transferred to our model. This enforcement of physical plausibility subsequently facilitates more effective pruning and regularization during the training process.
@@ VI Optimization Strategies for Cell Based Gaussian Splatting
The cell search rasterization inevitably introduces redundant 3D Gaussians during the blending stage, due to the discrepancies that arise between the spatial intervals of cells and the emitted neural Gaussians as shown in Fig. 6a. As the efficiency of cell search rasterization is contingent upon the degree of coherence between the visible regions of a cell and its associated Gaussians, we provide several engineering-driven strategies to facilitate training of Duplex-GS.
VI-A False Positive Skip
@@
Tile-based Rendering (TBR) is widely adopted in modern GPUs and is also utilized in 3DGS rasterization. However, this approach often introduces a substantial number of redundant false-positive Gaussians. This problem is further exacerbated in cell-based Gaussian frameworks due to the spatial mismatch between cell regions and the actual coverage of Gaussians. To mitigate the associated performance degradation, we introduce a pixel-level radius based culling strategy to eliminate unnecessary Gaussian kernel computations. Specifically, as illustrated in Fig. 6, a lightweight axis-aligned bounding box (AABB) check is employed: Gaussians are excluded from the blending operation if the pixel lies outside their respective AABBs. This efficient early-exit mechanism significantly reduces redundant computation and accelerates the overall rendering process.
VI-B Geometric Correction via Neural Gaussians
In addition to skipping false-positive Gaussians to improve efficiency, maximizing the overlap between each cell and its emitted neural Gaussians further reduces the number of redundant Gaussians. To achieve better geometric alignment, we dynamically define the cell’s quaternion as a weighted average of the quaternions associated with its emitted Gaussians, as described in Eq. 12. Moreover, to ensure an even spatial distribution of Gaussians within each cell, the cell centers are periodically updated using
@@ VI-B Geometric Correction via Neural Gaussians
VI-C Cell Position Reset
Following adaptive density control [11], we grow new cells at Gaussians with high view-space positional gradients. Unlike abstract anchors [20, 21], our cells have explicit geometry. To decouple existing cells from newly introduced ones, the offsets of the significant Gaussians are reset to zero after cell growth, i.e., their positions are reinitialized to the centers of the corresponding cells.
VI-D Perceptual Loss
The loss function for optimizing 3DGS is typically formulated as a combination of $\ell_{1}$ and SSIM [56] losses, which respectively capture pixel-level reconstruction error and structural similarity. As proven in GAN-based models, these pixel-wise image differences results in solutions with high signal-to-noise ratios, which are perceptually rather smooth and less convincing [57, 58]. We additionally incorporate a perceptual loss term for training Gaussian based models. The overall loss is defined as
@@ VII Experimental Results
    Experiments are conducted with our own implementation, which has been recognized by the author of [16]. Codes are available at https://github.com/LiYukeee/sort-free-gs.
  • 3
    Results from original paper (code unavailable). Missing data reflects source omissions.
@@
    Experiments are conducted with our own implementation, which has been recognized by the author of [16]. Codes are available at https://github.com/LiYukeee/sort-free-gs.
TABLE III: Quantitative comparisons on MatrixCity. Duplex-GS achieves highest accuracy across all metric in such large-scale aerial scenarios while maintaining competitive efficiency.
Method MatrixCity
$\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{\#GS/\#Proxy}_{\downarrow}$ $\text{FPS}_{\uparrow}$ $\text{Time~[h]}_{\downarrow}$
3DGS1 [11] 27.03 0.808 0.310 10206 K 66 2.60
Scaffold-GS [20] 26.82 0.811 0.288 4829 K 113 2.31
Octree-GS [21] 25.18 0.738 0.397 4082 K 128 2.61
Octree-GS* [21] 26.61 0.812 0.292 8010 K 79 -
LC-WSR [16] 26.34 0.782 0.340 7841 K 34 3.86
Ours $(K=5)$ 27.40 0.826 0.260 4165 K 119 2.21
    Officially released pretrained model.